OpenStudy (anonymous):

Solve: 2(x + 8) − (x + 1) = 20

OpenStudy (anonymous):

solve the brackets first

OpenStudy (anonymous):

u have to first to distribute property to 2(x+8) so it would equal 2x + 16 - (x + 1) =20

hero (hero):

Distribute 2 across x + 8 to get 2x + 16 Distribute -1 across x + 1 to get -x - 1 2x + 16 -x - 1 = 20 Place like terms on the same side: 2x - x = 20 + 1 - 16 Combine terms to get x = 5

OpenStudy (anonymous):

its easy just multiply and then subtract 20 on both sides !!